OpenStudy (anonymous):

What is the simplest form of the number? 16 ^3/4

OpenStudy (anonymous):

and this one WHAT IS THE solution to sqrt 2x-1=3

OpenStudy (anonymous):

\[16^\frac{3}{4} = 2^{4.\frac{3}{4}} = 2^3 = 8\] \[\sqrt{2x - 1} =3\] Then x = 5

OpenStudy (anonymous):

Thanks